# Evaluating Search-Based Software Microbenchmark Prioritization

### v2

@@ 1 Introduction
The results show that the best SBSMBP technique, i.e., the single-objective Generational Genetic Algorithm (GA) combining the three objectives with weighted-sum (GA C-CO-CH), performs competitively with the best greedy baseline, i.e., Total. However, it does not outperform Total regarding effectiveness, with both having an overall median APFD-P of $0.60$, while only adding a minor additional overhead of $1$% compared to below $1$% for Total. The best MOEA is Multi-Objective Cellular Genetic Algorithm (MOCell), which is not as effective as GA C-CO-CH and Total, with an overall median APFD-P of $0.58$. Surprisingly, the Greedy technique that relies only on the historical performance change size (i.e., Greedy CH) performs better than Total regarding median APFD-P of $0.65$. Statistically, the best Greedy and SBSMBP techniques perform equally to Total, regarding effectiveness. However, Greedy CH is more effective than Total for one of the studied projects and never worse, while having a significantly lower overhead of $1$% consistently across the studied projects, whereas the coverage-based techniques have overheads ranging between $8$% and $105$% across the projects and $17$% on average.
These results tell that the Greedy technique relying solely on the performance change history (CH) performs the best, and it is arguably the simplest to implement. Hence, we recommend practitioners to employ the Greedy CH technique to achieve consistently earlier performance change detection, e.g., as part of CI, and researchers to investigate non-coverage-based SMBP techniques in the future.
@@
  • three search objectives (two of which are novel) to use in SMBP algorithms;
  • an experimental study showing the effectiveness and efficiency of the new Greedy SMBP and SBSMBP techniques compared to greedy baselines;
@@ 2 Software Microbenchmarking with JMH
As measuring performance is non-deterministic and multiple factors can influence the measurement, one has to repeatedly execute each benchmark to get reliable results. JMH executes each combination of a benchmark method and its parameterization according to a specified configuration, set through annotations on class or method level (lines 11115555) or through the command line interface (CLI). Figure 1 visualizes the repeated execution of a benchmark suite. Each (parameterized) benchmark is invoked as often as possible for a defined time period (e.g., $1$s, configured on lines $2$ and $3$), called an iteration, and the measured performance metrics are reported. To get reliable results, JMH first runs a set of warmup iterations (line $2$ and “wi”) to get the system into a steady state, for which the measurements are discarded. After the warmup, JMH runs a set of measurement iterations (line $3$ and “mi”). To deal with non-determinism of the Java Virtual Machine (JVM), JMH repeats the sets of warmup and measurement iterations for a number of forks (line $1$ and “f”), each in a new JVM instance. The result of a benchmark is then the distribution of all measurement iterations (“mi”) of all forks (“f”). A microbenchmark suite usually contains many benchmarks, which JMH executes sequentially.
@@ 3 Search-Based Prioritization
The post-processing stage has two tasks. (1) Select one solution from the SA results. In the single-objective SA case there is only one solution. In the multi-objective SA case, the SA returns the Pareto front containing multiple solutions to select from [28]. (2) Adjust the benchmark order based on the source code changes, when using a change-awareness technique (see Section 4.3.4). Finally, SBSMBP executes the whole benchmark suite in the optimized order. That is, it executes each benchmark repeatedly according to performance engineering best practice [30], measures the performance and compares the measurements to the old version for change detection. The changes are then reported to the developers and stored for future versions, i.e., to compute the historical performance changes objective.
@@ 4.2.3 Coverage Information and SMBP Baselines
The dataset provides the necessary coverage information for the SBSMBP, Greedy SMBP, and greedy baseline techniques. Dynamic coverage information, i.e., method coverage, is extracted per benchmark and version with JaCoCo [^1] https://www.jacoco.org/jacoco/, by executing a benchmark once (with JMH’s single-shot mode) and injecting the JaCoCo agent. Based on these coverages, we [56] studied coverage-based, greedy SMBP techniques, particularly Total and Additional – our baselines.
@@ 4.3.3 Search Objectives
[…]
@@ 4.6 Statistical Analyses
We use the non-parametric Kruskal–Wallis H [49] test to compare multiple sets of observations. Note that Arcuri and Briand [6] suggest using the Mann-Whitney U test, however, which only compares two sets of observations. The Kruskal-Wallis H test is considered an extension of the Mann-Whitney U test and is, therefore, compatible with the best practice by Arcuri and Briand [6]. The null hypothesis H0 states that the different distributions’ medians are the same, and the alternative hypothesis H1 is that they are different. If H0 can be rejected, we apply Dunn’s post-hoc test [23] to identify which pairs of observations are statistically different.
@@ 5.1.1 RQ1a: Effectiveness of SBSMBP Techniques
From Table II’s “Overall” column, we observe that the median APFD-P ranges from $0.48$ to $0.60$. The best performing SBSMBP technique is GA C-CO-CH with $0.60$, followed by GA CH with $0.58$ and MOCell with $0.58$. Interestingly, GA CH performs competitively by only relying on the historical performance change size and without relying on coverage. We further notice that HC (all objectives) and GA CO exhibit the worst effectiveness.
@@
The single-objective GA C-CO-CH also performs the best statistically, better than all the MOEAs except IBEA with at least a small effect size. This is interesting because a simple (single-objective) GA using weighted-sum performs equally or better than the MOEAs, which shows that in our context, using a MOEA does not lead to higher effectiveness. In addition, using GA instead of any MOEA has the benefit that the output is a single benchmark ranking and not a Pareto front from which a solution has to be selected. The remaining GA and MOEAs strategies are statistically equivalent though they show a minor difference in median effectiveness.
@@
Table II shows the median APFD-P per project, and Fig. 4 depicts the pair-wise statistical test results. We observe that the effectiveness depends on the project the SMBP technique is applied to. E.g., for GA C-CO-CH, APFD-P ranges from $0.54$ (for Netty) to $0.71$ (for Xodus).
@@
RQ1a Summary: GA C-CO-CH is the most effective SBSMBP technique overall, as it performs the best for six of the ten projects. The other SBSMBPs techniques worth considering are GA CH, particularly because it does not require coverage information, and MOCell, which is the most effective MOEA.
@@ 5.1.2 RQ1b: Comparison to Greedy Techniques
A surprising result is that the SBSMBP techniques are either equal (GA C-CO-CH) or worse (GA CH and MOCell) than Total when considering the median effectiveness. However, our observation aligns with TCP research [63], where the search-based technique is inferior to the greedy baseline (in their case Additional). The statistical tests, depicted in Fig. 5, confirm that the SBSMBP techniques are statistically equal to but never worse than Total and more effective than Additional (with effect sizes small and medium).
The Greedy techniques show a similar pattern as the GA techniques in RQ1a: Greedy CO is the least effective technique overall with $0.44$ APFD-P, Greedy C is equivalent to Total at $0.60$, and both Greedy C-CO-CH with $0.61$ and Greedy CH with $0.65$ are more effective than Total, when considering the median. However, the improvement of the latter two Greedy techniques over Total is not statistically significant (see Fig. 5). Nevertheless, using Greedy CH and achieving a better median effectiveness even though being statistically equivalent is a surprising, yet great result, as it does not require coverage information. RQ2 in Section 5.2 investigates the implications of this on the technique efficiency.
@@
While overall the baseline Total is not statistically outperformed by any of the SBSMBP and Greedy techniques, the per-project results offer a nuanced view inTable III and Fig. 6.
@@
The two best Greedy techniques perform favorably compared to Total: (1) GreedyCHis better for one project (Byte Buddy) and equal for nine, while the medians are better for four projects, worse for five, and equal for one; (2) GreedyC-CO-CHis equal for all ten projects, while the medians are better for seven, worse for two, and equal for one.
@@
The most effective strategies according to the median for each project are: (1) GreedyCHfor four, i.e., Eclipse Collections, Log4j 2, RxJava, and Zipkin; (2) GreedyC-CO-CHfor three, i.e., JCTools, Netty, and Xodus; (3) Totalfor two, i.e., Jenetics and Log4j 2; (4) GAC-CO-CHfor two, i.e., Byte Buddy and JCTools; and (5) GACHfor one, i.e., Byte Buddy.
@@ 5.1.3 RQ1c: Impact of Change-Awareness
From the overall results, we observe differences among the change-aware approaches, although with no best approach across SMBP techniques and projects. Specifically, greedy techniques are less sensitive to change-awareness, i.e., change-awareness has hardly an impact on effectiveness. Whereas the SBSMBP techniques are considerably impacted by the change-awareness choice, each technique is not equally affected by the same approach. We notice that change-awareness is connected to the project that the SMBP is applied to, e.g., using cac for Byte Buddy is most effective, car for JCTools, and nca for (most SMBP techniques for) Okio. This shows that there often is a benefit for practitioners choosing the “right” change-awareness approach depending on the project and SMBP technique.
@@ 5.2 RQ2: Efficiency
While the previous section investigated the effectiveness of the SBSMBP techniques and Greedy techniques, a holistic evaluation requires an efficiency evaluation to understand whether using a technique is feasible in practice. This section compares the greedy heuristics Additional and Total, Greedy techniques, and SBSMBP techniques with respect to the overhead they impose on the overall benchmark suite execution time. Table IV shows the median and MAD efficiency overheads overall and per project. Overheads below $1$% are depicted as $<1$. We refrain from reporting statistical tests because they would arguably neither add valuable insights nor change the conclusions.
@@ C-CO-CH Prioritization $<11$$<11$$<11$$<11$$<11$$<11$$<11$$<11$$<11$$<11$$1$%$\pm$<$1$
Both baselines add less than $1$% for running the algorithm, i.e., prioritization time, and combined with the coverage extraction time have a total runtime overhead, i.e., analysis time, of $17$%. This shows that the analysis time is dominated by the coverage extraction and the SMBP algorithm plays only a minor role. These numbers are from the data set of our previous work [56]. The same applies to (most of) the Greedy techniques.
The single-objective SBSMBP techniques (i.e., HC and GA) have around $1$% prioritization time. This brings the analysis time to between $17$% and $19$%, similar to the greedy baselines. Greedy CH, HC CH, and GA CH, however, have a similar analysis and prioritization time. This is because the CH objective does not rely on coverage extraction, which takes the majority of the analysis time for coverage-based techniques. In particular, Greedy CH becomes (even more) appealing, as it is not only the most effective technique (or at least equal to Total) but also the most efficient one.
@@
The results per project show a diverse situation for the SMBP techniques that rely on coverage objectives: the analysis times range from between $8$% and $13$% for Jenetics to between $104$% and $105$% for Eclipse Collections. For eight of the projects, the analysis time is $23$% or below. For Eclipse Collections and Netty, where the overhead is $105$% and $49$% respectively, applying SMBP is not worthwhile.
These project-dependent overheads suggest that the techniques with non-coverage-based objectives are critical to be universally applicable across projects, especially when run as part of CI. Hence, the overhead results for the CH techniques are especially promising: all CH techniques have an analysis time overhead of $2$% or less across all the projects. When considering the most effective technique Greedy CH, the overhead even drops below $1$% for all the projects. Consequently, the analysis time of the techniques with non-coverage-based objectives is solely dependent on the prioritization time (i.e., the time taken to run the SMBP algorithm), which will likely always be considerably smaller than the extensive benchmark suite runtimes; therefore, these SMBP techniques are generally beneficial being applied to microbenchmarks.
@@ 6 Discussion
The results show that the best SBSMBP technique (GA C-CO-CH) is only competitive with the best greedy baseline (Total). When using a simple Greedy technique that only considers the historical performance change size (CH), one can retrieve more effective SMBP rankings with significantly lower overhead than when relying on coverage objectives (C and CO), which is the case for both the greedy baselines and the SBSMBP techniques. This section discusses our empirical results from several different perspectives.
@@
First, coverage is not “good enough” as objectives for SMBP. This becomes evident from this and our previous study [56]. Furthermore, Chen et al. [11] observed that in the context of benchmark selection, functional bug prediction metrics are more important than code-level performance metrics. This suggests that future SMBP techniques should explore dedicated metrics related to size, diffusion, or history as well as specific code-level performance metrics, such as loops or synchronization (similar to Laaber et al. [55]). However, given the high coverage extraction overhead, non-code metrics should be more favoured than code ones.
Second, optimal hyperparameters of the SAs could improve SBSMBP effectiveness. However, we explored the effectiveness of the SAs with more generations and investigated the convergence of the objectives and found that the objectives converge quickly. This suggests that the search space is relatively trivial (for the given objectives) and SAs do not offer an advantage over simple greedy techniques, which our results empirically show. Moreover, the search objectives we employed, in particular the coverage-based objectives, are not effective in finding bigger performance changes sooner, and other objectives should be explored in the future (such as the ones mentioned above).
@@
Is it Practical to Apply SMBP? Beyond the temporal cost of applying a SMBP technique, there is a cost of extracting and maintaining the information required as input. On the one hand, the techniques require storing historical information about benchmark executions, especially their changes between previous versions, for computing the CH objective. On the other hand, there is neither a need to store historical coverage information nor compute the source code difference on every new commit, as the results of RQ1c show that nca, the non-change-aware approach, perform equal to the change-aware approaches (i.e., cac and car) overall. This is probably best done as part of the CI pipeline because it provides the necessary infrastructure to run benchmarks, store the information about the build (i.e., performance change history), and has the code changes readily available.
@@
Research on functional TCP often uses datasets with seeded, artificial faults (mutations) to evaluate TCP effectiveness (e.g., [64, 22]). Conversely, performance regression testing research mostly relies on measured performance changes between adjacent versions [67, 56]. This is due to performance mutation testing having only recently received attention [20, 42] and large datasets, such as Defects4J [44], not existing. Because of the enormous experiment runtimes required for rigorously executing benchmarks (this study’s dataset took $89$ days to create), performing performance mutation testing, where for each mutant the whole benchmark suite has to be executed, becomes unrealistic. However, this does not invalidate the findings of this study or any other experimental performance regression study relying on measured performance changes as long as the measurements have been conducted rigorously.
@@
What is an Important Performance Change? Following previous research on SMBP [67, 56], this paper also defines the importance of a benchmark as the performance change size it detects. A benchmark that detects a large performance change is considered more important than a benchmark that detects a small one. Accordingly, we conclude whether a technique is more or less effective. It is, however, unclear whether this definition of importance is “accurate.” Alternative formulations of importance could be based on: (1) the impact of the code called by the benchmark on the performance of an application, (2) the project developer’s perception and context-dependent knowledge, or (3) the usage frequency of the code from application programming interface (API) clients. Moreover, developers might consider a performance change, irrespective of its size, as important. All these different definitions of importance would likely change this paper’s results and require a study of its own.

### v4

@@ 1 Introduction
The results show that the best SBSMBP technique, i.e., the single-objective Generational Genetic Algorithm (GA) combining the three objectives with weighted-sum (GA C-CO-CH), performs competitively with the best greedy baseline, i.e., Total. However, it does not outperform Total regarding effectiveness, with both having an overall median APFD-P of $0.60$, while only adding a minor additional overhead of $1$% compared to below $1$% for Total on top of the coverage extraction overhead. The best MOEA is Multi-Objective Cellular Genetic Algorithm (MOCell), which exhibits a lower median APFD-P of $0.58$ than GA C-CO-CH and Total. Surprisingly, the Greedy technique that relies only on the historical performance change size (i.e., Greedy CH) has a higher median APFD-P of $0.65$ than Total. Statistically, however, the best Greedy and SBSMBP techniques are only as effective as Total. Per project, Greedy CH is more effective than Total for one project and never worse, while having a significantly lower overhead of $1$% consistently across the studied projects, whereas the coverage-based techniques have overheads ranging between $8$% and $105$% across the projects and $17$% on average.
These results reveal that the Greedy technique relying solely on the performance change history (CH) performs the best and is arguably the simplest to implement. Hence, we recommend practitioners to employ the Greedy CH technique to achieve consistently earlier performance change detection, e.g., as part of CI, and researchers to investigate non-coverage-based SMBP techniques in the future.
@@
  • three search objectives (two of which are novel) to employ in SMBP algorithms;
  • an experimental study showing the effectiveness and efficiency of the new Greedy SMBP and SBSMBP techniques compared to greedy, coverage-based baselines;
@@ 2 Software Microbenchmarking with JMH
[…]
@@ 3 Search-Based Prioritization
The post-processing stage has two tasks. (1) Select one solution from the SA results. In the single-objective SA case there is only one solution. In the multi-objective SA case, the SA returns the Pareto front containing multiple solutions to select from [28]. (2) Adjust the benchmark order based on the source code changes, when using a change-aware technique (see Section 4.3.4). Finally, SBSMBP executes the whole benchmark suite in the optimized order. That is, it executes each benchmark repeatedly according to performance engineering best practice [30], measures the performance, and compares the measurements to the old version for change detection. The changes are then reported to the developers and stored for future versions, i.e., to compute the historical performance changes objective.
@@ 4.2.3 Coverage Information and SMBP Baselines
The dataset provides the necessary coverage information for the SBSMBP, Greedy SMBP, and greedy baseline techniques. Dynamic coverage information, i.e., method coverage, is extracted per benchmark and version with JaCoCo [^1] https://www.jacoco.org/jacoco/, by executing a benchmark once (with JMH’s single-shot mode) and injecting the JaCoCo agent. Based on these coverages, we [56] studied coverage-based, greedy SMBP techniques, particularly Total and Additional — our baselines.
@@ 4.3.3 Search Objectives
$\textit{APTEC}=\frac{\sum\limits^{n}_{b=1}\frac{prev(b-1)+elements(b)}{m_{total}}}{n}$ (1)
@@ 4.6 Statistical Analyses
We use the non-parametric Kruskal–Wallis H [49] test to compare multiple sets of observations. Note that Arcuri and Briand [6] suggest using the Mann-Whitney U test, which, however, only compares two sets of observations. The Kruskal-Wallis H test is considered an extension of the Mann-Whitney U test and is, therefore, compatible with the best practice by Arcuri and Briand [6]. The null hypothesis H0 states that the different distributions’ medians are the same, and the alternative hypothesis H1 is that they are different. If H0 can be rejected, we apply Dunn’s post-hoc test [23] to identify which pairs of observations are statistically different.
@@ 5.1.1 RQ1a: Effectiveness of SBSMBP Techniques
From Table II’s “Overall” column, we observe that the median APFD-P ranges from $0.48$ to $0.60$. The SBSMBP technique with the highest median APFD-P is GA C-CO-CH with $0.60$, followed by GA CH with $0.58$ and MOCell with $0.58$. Interestingly, GA CH performs competitively by only relying on the historical performance change size and without relying on coverage. We further notice that HC (all objectives) and GA CO exhibit the worst effectiveness.
@@
The single-objective GA C-CO-CH also performs the best statistically, better than all the other SAs except GA CH and IBEA with at least a small effect size. This is interesting because a simple (single-objective) GA using weighted-sum performs equally or better than the MOEAs, which shows that in our context, using a MOEA does not lead to higher effectiveness. In addition, using GA instead of any MOEA has the benefit that the output is a single benchmark ranking and not a Pareto front from which a solution has to be selected. The remaining GA and MOEA strategies are statistically equivalent though they show a minor difference in median effectiveness.
@@
Table II shows the median APFD-P per project, and Fig. 4 depicts the pair-wise statistical test results. We observe that the effectiveness depends on the project the SMBP technique is applied to, e.g., for GA C-CO-CH, APFD-P ranges from $0.54$ (for Netty) to $0.71$ (for Xodus).
@@
RQ1a Summary: GA C-CO-CH is the most effective SBSMBP technique overall, as it performs the best for six of the ten projects. The other SBSMBP techniques worth considering are GA CH, particularly because it does not require coverage information, and MOCell, which is the most effective MOEA.
@@ 5.1.2 RQ1b: Comparison to Greedy Techniques
A surprising result is that the SBSMBP techniques exhibit either only an equal (GA C-CO-CH) or a lower (GA CH and MOCell) median APFD-P compared to Total. However, our observation aligns with TCP research [63], where the search-based technique is inferior to the greedy baseline (in their case Additional). The statistical tests, depicted in Fig. 5, confirm that the SBSMBP techniques are statistically equal to but never worse than Total and better than Additional (with small and medium effect sizes).
The Greedy techniques show a similar pattern as the GA techniques in RQ1a: Greedy CO is the technique with the lowest median APFD-P of $0.44$, Greedy C is equivalent to Total at $0.60$, and both Greedy C-CO-CH with $0.61$ and Greedy CH with $0.65$ exhibit higher median APFD-Ps than Total. However, the improvement of the latter two Greedy techniques over Total is not statistically significant (see Fig. 5). Nevertheless, using Greedy CH and achieving a better median effectiveness even though being statistically equivalent is a surprising, yet great result, as it does not require coverage information. RQ2 in Section 5.2 investigates the implications of this on the technique efficiency.
@@
While overall the baseline Total is not statistically outperformed by any of the SBSMBP and Greedy techniques, the per-project results offer a nuanced view in Tables III and 6.
@@
The two best Greedy techniques perform favorably compared to Total: (1) GreedyCHis better for one project (Byte Buddy) and equal for nine, while the medians are better for four projects, worse for five, and equal for one; (2) GreedyC-CO-CHis equal for all the ten projects, while the medians are better for seven, worse for two, and equal for one.
@@
The strategies with the highest median APFD-P for each project are: (1) GreedyCHfor four, i.e., Eclipse Collections, Log4j 2, RxJava, and Zipkin; (2) GreedyC-CO-CHfor three, i.e., JCTools, Netty, and Xodus; (3) Totalfor two, i.e., Jenetics and Log4j 2; (4) GAC-CO-CHfor two, i.e., Byte Buddy and JCTools; and (5) GACHfor one, i.e., Byte Buddy.
@@ 5.1.3 RQ1c: Impact of Change-Awareness
From the overall results, we observe differences among the change-aware approaches, although with no best approach across SMBP techniques and projects. Specifically, greedy techniques are less sensitive to change-awareness, i.e., change-awareness has hardly an impact on the effectiveness. Whereas the SBSMBP techniques are considerably impacted by the change-awareness choice, each technique is not equally affected by the same approach. We notice that change-awareness is connected to the project that the SMBP is applied to, e.g., using cac for Byte Buddy, car for JCTools, and nca for (most SMBP techniques for) Okio is most effective. This shows that there often is a benefit for practitioners choosing the “right” change-awareness approach depending on the project and SMBP technique.
@@ 5.2 RQ2: Efficiency
While the previous section investigated the effectiveness of the SBSMBP and Greedy techniques, a holistic evaluation requires an efficiency evaluation to understand whether using a technique is feasible in practice. This section compares the greedy heuristics Additional and Total, Greedy techniques, and SBSMBP techniques with respect to the overhead they impose on the overall benchmark suite execution time. Table IV shows the median and MAD efficiency overheads overall and per project. Overheads below $1$% are depicted as $<1\%$. We refrain from reporting statistical tests, as they would arguably neither add valuable insights nor change the conclusions.
@@ C-CO-CH Prioritization $<11$$<11$$<11$$<11$$<11$$<11$$<11$$<11$$<11$$<11$$1$%$\pm$<$1$
Both baselines add less than $1$% for running the algorithm, i.e., prioritization time, and combined with the coverage extraction time have a total runtime overhead, i.e., analysis time, of $17$%. This shows that the analysis time is dominated by the coverage extraction and the SMBP algorithm plays only a minor role. These numbers are from our previous work [56]. The same applies to the coverage-based Greedy techniques, i.e., C, CO, and C-CO-CH.
The single-objective SBSMBP techniques (i.e., HC and GA) have around $1$% prioritization time. This brings the analysis time to between $17$% and $19$%, similar to the greedy baselines. Greedy CH, HC CH, and GA CH, however, have a similar analysis and prioritization time. This is because the CH objective does not rely on coverage extraction, which takes the majority of the analysis time for coverage-based techniques. In particular, Greedy CH becomes (even more) appealing, as it is not only among the most effective techniques (with, e.g., Total) but also the most efficient one.
@@
The results per project show a diverse situation for the SMBP techniques that rely on coverage objectives: the analysis times range from between $8$% and $13$% for Jenetics to between $104$% and $105$% for Eclipse Collections. For eight of the projects, the analysis time is $23$% or less. For Eclipse Collections and Netty, where the overhead is $105$% and $49$% respectively, applying SMBP is not worthwhile.
These project-dependent overheads suggest that the techniques with non-coverage-based objectives are critical to be universally applicable across projects, especially when run as part of CI. Hence, the overhead results for the CH techniques are especially promising: all CH techniques have an analysis time overhead of $2$% or less across all the projects. When considering the most effective technique Greedy CH, the overhead even drops below $1$% for all the projects. Consequently, the analysis time of the techniques with non-coverage-based objectives is solely dependent on the prioritization time (i.e., the time taken to run the SMBP algorithm), which will likely always be considerably smaller than the extensive benchmark suite runtimes; therefore, it is generally beneficial to apply these SMBP techniques.
@@ 6 Discussion
The results show that the best SBSMBP technique (GA C-CO-CH) is only competitive with the best greedy baseline (Total). When using a simple Greedy technique that only considers the historical performance change size (CH), one can retrieve equally and sometimes more effective SMBP rankings with significantly lower overhead than when relying on coverage objectives (C and CO), which is the case for both the greedy baselines and the SBSMBP techniques. This section discusses our empirical results from several different perspectives.
@@
First, coverage is not “good enough” as objectives for SMBP. This becomes evident from this and our previous study [56]. Furthermore, Chen et al. [11] observed that in the context of benchmark selection, functional bug prediction metrics are more important than code-level performance metrics. This suggests that future SMBP techniques should explore dedicated metrics related to size, diffusion, or history as well as specific code-level performance metrics, such as loops or synchronization (similar to Laaber et al. [55]). However, given the high coverage extraction overhead, non-code metrics should be more favoured over code metrics.
Second, optimal hyperparameters of the SAs could improve SBSMBP effectiveness. For this, we explored the SAs’ effectiveness with more generations, investigated the convergence of the objectives, and found that the objectives converge quickly. This suggests that the search space is relatively trivial (for the given objectives) and SAs do not offer an advantage over simple greedy techniques, which our results empirically show. Moreover, the employed search objectives, in particular the coverage-based objectives, are ineffective in finding larger performance changes sooner, and other objectives should be explored in the future (such as the ones mentioned above).
@@
Is it Practical to Apply SMBP? Beyond the temporal cost of applying a SMBP technique, there is a cost of extracting and maintaining the information required as input. On the one hand, the techniques require storing historical information about benchmark executions, especially their changes between previous versions, for computing the CH objective. On the other hand, there is neither a need to store historical coverage information nor compute the source code difference on every new commit, as the results of RQ1c show that nca, the non-change-aware approach, performs equally to the change-aware approaches (i.e., cac and car) overall. This is probably best done as part of the CI pipeline because it provides the necessary infrastructure to run benchmarks, store the information about the build (i.e., performance change history), and has the code changes readily available.
@@
Research on functional TCP often uses datasets with seeded, artificial faults (mutations) to evaluate TCP effectiveness (e.g., [64, 22]). Conversely, performance regression testing research mostly relies on measured performance changes between adjacent versions [67, 56]. This is due to performance mutation testing having only recently received attention [20, 42] and large datasets, such as Defects4J [44], do not exist. Because of the enormous experiment runtimes required for rigorously executing benchmarks (this study’s dataset took $89$ days to create), performing performance mutation testing, where for each mutant the whole benchmark suite has to be executed, becomes unrealistic. However, this does not invalidate the findings of this study or any other experimental performance regression study relying on measured performance changes as long as the measurements have been conducted rigorously.
@@
What is an Important Performance Change? Following previous research on SMBP [67, 56], this paper also defines the importance of a benchmark as the performance change size it detects. A benchmark that detects a large performance change is considered more important than a benchmark that detects a small one. Accordingly, we conclude whether a technique is more or less effective. It is, however, unclear whether this definition of importance is “accurate.” Alternative formulations of importance could be based on: (1) the impact of the code called by the benchmark on the performance of an application, (2) the project developer’s perception and context-dependent knowledge, or (3) the usage frequency of the code from application programming interface (API) clients. Moreover, developers might consider any performance change as important, irrespective of the size. All these different definitions of importance would likely change our results and require a study of its own.